# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v1

@@ 5.3 Task-Agnostic Context for Feature Compression
The leftover spatial dependencies of $\boldsymbol{\mathrm{y}}=g_{a}(\boldsymbol{\mathrm{x}};\theta)$ are commonly around high-contrast areas. A compression model can adaptively (i.e., for each sample) capture leftover dependencies by including side information with a hyperprior [34, 28].
@@
Consider that high-contrast areas typically correspond to edges and other regions of interest, i.e., they commonly correspond to keypoints. Figure 7 contrasts leftover pixel dependencies of $\hat{y}$ from a powerful LIC model [28] to keypoints output by a KeyNet [55] network. Therefore, we should further improve compression performance with side information by embedding keypoints as context for encoding as follows:
@@
where $k$ is a keypoint extraction function $k:\mathbb{R}^{3\times H\times W}\rightarrow\mathbb{R}^{1\times H\times W}$ , $f_{ds}$ is a parametric downsampling function $f_{ds}:\mathbb{R}^{1\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times\frac{H}{2^{n}}\times\frac{W}{2^{n}}}$, and $a_{c}$ is a single (2D) cross-attention block (Figure 6).
The cross-attention block takes context as an additional input for weighting the latent with attention scores. For $k$, we use scores from a (frozen) pre-trained and simplified KeyNet [55] due to its robustness in diverse environments and low memory requirements (less than $6000$ parameters). While this method should generalize to LIC, it complements feature compression exceptionally well. We found quantizing and compressing $\boldsymbol{\mathrm{y}}_{ca}$ (i.e., $\boldsymbol{\tilde{\mathrm{h}}}=g_{s}(Q(\boldsymbol{\mathrm{y}}_{ca}),\psi)$ further lowers the bitrate without affecting task performance.
@@ 5.4 Compression Model Architecture
For the remainder of the work, we refer to the compression model as an encoder-decoder pair $enc,dec$. The encoder comprises $g_{a},h_{a},k,f_{ds},c_{a}$, and the entropy coder. The decoder consists of the $g_{s}$ and the entropy decoder. The entropy model $p_{\boldsymbol{\hat{\mathrm{z}}}}(\boldsymbol{\hat{\mathrm{z}}})$ and $h_{s}$ are shared.
@@ 5.5 Single Encoder with Multiple Backbones and Tasks
Analogous to FrankenSplit [17], consider a set of $n$ shallow and deep layers pairs of backbones (i.e., foundational models):
@@ 6 Evaluation
6.1 Experiment Setting and Methodology
@@ 6.1.1 Testbed
Device
CPU
GPU
Ground Server
16x Ryzen @ 3.4 GHz
RTX 4090
Edge (Nano Orin)
6x Cortex @ 1.5 GHz
Amp. 512 CC 16 TC
Edge (TX2)
4x Cortex @ 2 GHz
Pas. 256 CC
Edge (Xavier NX)
4x Cortex @ 2 GHz
Vol. 384 CC 48 TC
Our simulation replicates a configurable CubeSat by imposing constraints on energy, memory, and bandwidth. Table II summarizes the missions [6, 60, 61, 62] and links we use experiment traces on.
@@
Oper. Constellation Link
Rate (Mbps)
Pass (s)
Data Per Pass (GB)
Planet Dove (3P/B13) HSD 1
160
510
12.0
Maxar WorldView WorldView-3
1200
600
90
ESA Copernicus Sentinel 3A/B
560
600
40.0
NASA Landsat Landsat 8
440
120
39.6
@@ 6.1.3 Training & Implementation Details
We fix the tile resolution to $512\times 512$ during training. We load samples as a video sequence for FOOL by grouping tiles from the same image in partitioning order with random transformations to fill any remaining spots. After training, the tensor shape (i.e., the number of tiles and the spatial dimensions) may vary for each separate sample.
We use PyTorch [75], CompressAI [76], and pre-trained detection models from Ultralytics [63]. To ensure reproducibility, we use torchdistill [77].
We use an Adam optimizer [78] with a batch size of $8$ and start with an initial learning rate of $1\cdot 10^{-3}$, then gradually lower it to $1\cdot 10^{-6}$ with an exponential scheduler. We first seek a weight for the rate term in Equation 15 that results in lossless prediction with the lowest (best) bpp. Then, we progressively increase the term weight to evaluate trade-offs between rate and predictive loss.
@@ 6.1.6 Baselines
Codec
Pars. Total
Pars. Enc.
Pars. Dec.
Shared
FOOL-L
4.97M
1.19M
4.06M
0.28M
FOOL-M
4.33M
0.69M
3.83M
0.16M
FOOL-S
3.91M
0.35M
3.66M
0.08M
SVBI-L
4.99M
1.25M
4.39M
0.65M
SVBI-M
4.35M
0.72M
3.91M
0.29M
SVBI-S
3.95M
0.38M
3.71M
0.16M
FP
7.03M
3.51M
3.51M
0.02M
MSHP
17.56M
14.06M
11.66M
8.15M
JAHP
25.50M
21.99M
19.60M
16.10M
TinyLIC
28.34M
21.23M
19.16M
12.05M
@@ 6.2 Rate Trade-off with Predictive Loss
We report predictive loss with mAP@50 since the absolute differences to lossless prediction are more apparent than mAP@[50:95] between the decreasing quality levels.
@@ 6.2.1 Comparison to Image Codecs
We primarily focus on how FOOL compares to existing SVBI to draw new insights from the novel additions. The work in [17, 15, 16], already extensively demonstrate the advantage of SVBI over LIC models empirically when the objective is to minimize predictive loss for arbitrary tasks associated with a wide range of popular feature extractor architectures. We confirm that our results on aerial and satellite imagery datasets with one-stage detectors are consistent with previous findings.
@@ 6.2.3 Ablation Study
Model
DOTA-2 (bpp ↓↓)
SpaceNet-3 (bpp ↓↓)
xView (bpp ↓↓)
FOOL-L
0.1843
0.1760
0.1822
FOOL-M
0.2110
0.1993
0.2032
FOOL-S
0.2389
0.223
0.2290
BSVBI-L
0.3622
0.3440
0.3452
BSVBI-M
0.3775
0.3605
0.3699
BSVBI-S
0.3889
0.3852
0.3909
NAIT-L
0.2209
0.1993
0.2032
NAIT-M
0.2287
0.2193
0.2205
NAIT-S
0.2433
0.2386
0.2407
NKPC-L
0.2839
0.2712
0.2768
NKPC-M
0.2926
0.2855
0.2883
NKPC-S
0.3116
0.3029
0.3112
@@
The results show that relative to BSVBI, the task-agnostic context component contributes considerably more to rate reductions than the AIT layers that leverage inter-tile spatial dependencies. Still, we argue that the NTI-layers fulfill their purpose, to synergize with the partitioning strategy that maximizes processing throughput (Section 6.4).
6.3 Reconstruction for Human Perception
We aim to demonstrate the feasibility of recovering presentable images from the compressed latent space of shallow features. We average results on DOTA-2, SpaceNet-3, and xView to reduce the bloat of reporting similar values. For transparency, we exclusively select samples from the lower quartile across all measures to qualitatively showcase the reconstruction.
@@ 6.3 Reconstruction for Human Perception
Model
PSNR↑↑
MS-SSIM↑↑
LPIPS↓↓
BPP↓↓
Pred. Loss↓↓
FOOL
36.51
15.43
0.1480
0.1808
FOOL-FT 35.56 14.57 0.7100 0.1808 -
FP-HQ
43.22
25.07
0.0896
1.0470
1.500
@@
MSHP-HQ
43.90
25.20
0.0841
1.0370
1.711
MSHP-MQ 36.45 16.83 0.2361 0.2787 7.180
JAHP-HQ
43.95
25.17
0.0818
1.0297
1.504
@@
Figure 13 further visualizes the potential of reliably recovering fine-grained from the compressed latent space. Naturally, it should be possible to finetune the TinyLic-MQ to improve perceptual quality analogous to FOOl-FT. However, TinyLIC is still a significantly costlier model, with a worse rate and prediction performance.
More pressingly, we stress that the reliability of a restoration model is bound by the available signals in the compressed latent space. Accordingly, we deliberately avoid generative models prone to hallucination (e.g., Generative adversarial networks) that may inflate the significance of our work with more visually pleasing images. In particular, prioritizing perceptual quality over reliability defeats the primary purpose of image restoration, i.e., intervention by human experts in critical EO applications. A model outperforming experts does not imply that predictions may inexplicably be false. Specifically, where human cost is involved (e.g., disaster warning or relief [5]) it is paramount that experts can trust the codec to not include extrapolated elements to an image.
We argue that our results adequately underpin the statements in Section 3.3.1 and Section 3.3.2. In summary, if the salient regions align, compressing for machine perception (i.e., model prediction) requires more information than for human perception. Task specificity determines rate savings and not an entity’s input interface. Targeting shallow features is minimally task-specific by relaxing the objective for lossless prediction on all possible tasks to those valuable for clients.
@@ 6.4.1 Processing Throughput and Transfer Cost Reduction
Device Model
Spatial Dim.
Batch Size
TCR/s ↑↑
FOOL-L
600x600
4
$7.4794\cdot 10^{8}$
@@
Orin Nano FOOL-S
600x600
8
$\boldsymbol{1.3386\cdot 10^{9}}$
FOOL-L
600x600
8
$7.5456\cdot 10^{8}$
@@
NX FOOL-S
600x600
8
$\boldsymbol{1.3422\cdot 10^{9}}$
@@
Orin Nano BSVBI-S
550x550
7
$6.9919\cdot 10^{9}$
@@ 6.4.2 Model Inference with Concurrent Task Execution
Model
TCR/s [[[[conc]]]]
TCR/s dec.
File Size (MB)
File/s
rANS conc. (MB/s)
FOOL-L
$7.26\cdot 10^{8}$
2.94%
0.616
29
37.2
FOOL-M
$7.57\cdot 10^{8}$
1.28%
0.462
31
38.3
FOOL-S
$1.32\cdot 10^{9}$
1.06%
0.383
52
38.8
BSVBI-L
$5.96\cdot 10^{8}$
2.88%
0.822
37
37.6
BSVBI-M
$7.15\cdot 10^{8}$
1.37%
0.617
42
39.6
BSVBI-S
$6.91\cdot 10^{8}$
1.28%
0.437
58
40.1
@@
Figure 15: CPU Usage and GPU Usage of Encoder Network
@@
Figure 16: CPU and GPU Usage of Concurrent Pipeline
@@ 6.4.3 Downlinkable Data Volume
Figure 17: Downlinked Deecompressed Data Volume
Notice the logarithmic scale, i.e., FOOL improves downlinking using bent pipes by over two orders of magnitude without relying on prior information on the downstream tasks or crude filtering methods. For example, given Maxar’s WorldView-3 conditions [60], it would be possible to downlink roughly 9TB of sensor data per pass, before reaching downlink saturation. As a comparison, the state-of-the-art filtering method in [27] reports a 3$\times$ improvement, based on a definition of value.
Note that to provide a realistic presentation of the opportunities SVBI provides, we assume that a nanosatellite processes tiles until reaching a downlink segment. Moreover, we disregard the “computational deadline”, i.e., it can process all the data before reaching a ground segment. This is reasonable since there should always be enough data to process. If not produced by a single onboard sensor, constellations may designate certain satellites as compression nodes using reliable, high-capacity local communication channels [80]. Further, from the throughput evaluation and the orbital period, it is inferrable that even the low-end current-generation Orin Nano would barely miss the computational deadline. We remind the reader that the presented throughput measures represent a worst-case lower bound as no vendor-specific software or hardware optimization was applied.
@@ 6.4.4 Energy Consumption and Savings
Figure 18: Energy Cost of Compression Pipeline
As expected, processing on TX2 requires more energy than on NX and Orin Nano as it is slower. Somewhat interesting is that the NX consumes more energy than the Orin Nano. As they execute the same models with comparable processing throughput, the results suggest that the newer Jetson lineup is more energy-efficient.
Lastly, we measure savings from reduced transmission time, arguably an often undervalued advantage of compression. Admittedly, satellites will downlink as bandwidth permits, i.e., the transmission energy cost does not depend on the codec performance when there is saturation. Nonetheless, to intuitively show the amount of energy large volumes might require, we contrast with bent pipes in Figure 19.
@@ 7 Discussions and Limitations
The following elaborates on non-negligible limitations and highlights some research directions we find particularly promising to increase mission value. A general recommendation is to leverage recent advancements in image restoration [81] and utilize downlinked information to minimize the risk of hallucination.

### v2

@@ 5.3 Task-Agnostic Context for Feature Compression
The leftover spatial dependencies after encoding are commonly around high-contrast areas. Consider that high-contrast areas typically correspond to edges and other regions of interest, i.e., keypoints.
@@
As an example, Figure 7 contrasts leftover pixel dependencies of $\hat{y}$ from a LIC model [35] to keypoints output by a KeyNet [54] network. Therefore, we should further improve compression performance with side information by embedding keypoints as context for encoding as follows:
@@
where $k$ is a keypoint extraction function $k:\mathbb{R}^{3\times H\times W}\rightarrow\mathbb{R}^{1\times H\times W}$ , $f_{ds}$ is a parametric downsampling function $f_{ds}:\mathbb{R}^{1\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times\frac{H}{2^{n}}\times\frac{W}{2^{n}}}$, and $a_{c}$ is a single (2D) cross-attention block (Figure 6). The cross-attention block takes context as an additional input for weighting the latent with attention scores. For $k$, we use scores from a (frozen) pre-trained and simplified KeyNet [54] due to its robustness in diverse environments and low memory requirements (less than $6000$ parameters). While this method should generalize to LIC, it complements feature compression exceptionally well and found quantizing and compressing $\boldsymbol{\mathrm{y}}_{ca}$ (i.e., $\boldsymbol{\tilde{\mathrm{h}}}=g_{s}(Q(\boldsymbol{\mathrm{y}}_{ca}),\psi)$ further lowers the bitrate without affecting task performance.
@@ 5.4 Compression Model Architecture
For the remainder of the work, we refer to the compression model as an encoder-decoder pair $enc,dec$. The encoder comprises $g_{a},h_{a},k,f_{ds},c_{a}$, and the entropy coder. The decoder consists of the $g_{s}$ and the entropy decoder. The entropy model $p_{\boldsymbol{\hat{\mathrm{z}}}}(\boldsymbol{\hat{\mathrm{z}}})$ and $h_{s}$ are shared. Note that, despite deploying more components on the constrained sender, the encoder has significantly fewer parameters than the decoder, since we increase the width of the receiver-exclusive components.
@@ 5.5 Single Encoder with Multiple Backbones and Tasks
Analogous to [17], consider a set of $n$ shallow and deep layers pairs of backbones (i.e., foundational models):
@@ 6 Evaluation
6.1 Experiment Design and Methodology
@@ 6.1.1 Testbed
Device CPU GPU
Ground Server 16x Ryzen @ 3.4 GHz RTX 4090
Edge (Nano Orin) 6x Cortex @ 1.5 GHz Amp. 512 CC 16 TC
Edge (TX2) 4x Cortex @ 2 GHz Pas. 256 CC
Edge (Xavier NX) 4x Cortex @ 2 GHz Vol. 384 CC 48 TC
Our simulation replicates a configurable CubeSat by imposing energy, memory, and bandwidth constraints. To simulate the downlink bottleneck with varying link conditions, parameterize link conditions and data volume (Section 3.1) using real-world missions [6, 29, 59, 60] as summarized in Table II.
@@
Oper. Constellation Link Rate (Mbps) Pass (s) Data Per Pass (GB)
Planet Dove (3P/B13) HSD 1 160 510 12.0
Maxar WorldView WorldView-3 1200 600 90
ESA Copernicus Sentinel 3A/B 560 600 40.0
NASA Landsat Landsat 8 440 120 39.6
@@ 6.1.3 Training & Implementation Details
We fix the tile resolution to $512\times 512$ during training. We load samples as a video sequence for FOOL by grouping tiles from the same image in partitioning order with random transformations to fill any remaining spots. After training, the tensor shape (i.e., the number of tiles and the spatial dimensions) may vary for each separate sample. We use PyTorch [73], CompressAI [74], and pre-trained detection models from Ultralytics [61]. To ensure reproducibility, we use torchdistill [75]. We use an Adam optimizer [76] with a batch size of $8$ and start with an initial learning rate of $1\cdot 10^{-3}$, then gradually lower it to $1\cdot 10^{-6}$ with an exponential scheduler. We first seek a weight for the rate term in Equation 19 that results in lossless prediction with the lowest (best) bpp. Then, we progressively increase the term weight to evaluate trade-offs between rate and predictive loss.
@@ 6.1.6 Baselines
Codec Pars. Total Pars. Enc. Pars. Dec. Shared
FOOL-L 4.97M 1.19M 4.06M 0.28M
FOOL-M 4.33M 0.69M 3.83M 0.16M
FOOL-S 3.91M 0.35M 3.66M 0.08M
SVBI-L 4.99M 1.25M 4.39M 0.65M
SVBI-M 4.35M 0.72M 3.91M 0.29M
SVBI-S 3.95M 0.38M 3.71M 0.16M
FP 7.03M 3.51M 3.51M 0.02M
MSHP 17.56M 14.06M 11.66M 8.15M
JAHP 25.50M 21.99M 19.60M 16.10M
TinyLIC 28.34M 21.23M 19.16M 12.05M
@@ 6.2 Rate Trade-off with Predictive Loss
We report the predictive loss as a percentage point difference using mAP@50 on foundational detection models.
@@ 6.2.1 Comparison to Image Codecs
We primarily focus on how FOOL compares to existing SVBI to draw new insights from the novel additions and confirm that our results on aerial and satellite imagery datasets with one-stage detectors are consistent with previous findings on standardized terrestrial datasets  [17, 15, 16].
@@ 6.2.3 Ablation Study
Model DOTA-2 (bpp ↓↓) SpaceNet-3 (bpp ↓↓) xView (bpp ↓↓)
FOOL-L 0.1843 0.1760 0.1822
FOOL-M 0.2110 0.1993 0.2032
FOOL-S 0.2389 0.223 0.2290
BSVBI-L 0.3622 0.3440 0.3452
BSVBI-M 0.3775 0.3605 0.3699
BSVBI-S 0.3889 0.3852 0.3909
NITA-L 0.2209 0.1993 0.2032
NITA-M 0.2287 0.2193 0.2205
NITA-S 0.2433 0.2386 0.2407
NKPC-L 0.2839 0.2712 0.2768
NKPC-M 0.2926 0.2855 0.2883
NKPC-S 0.3116 0.3029 0.3112
@@
The results show that relative to BSVBI, the task-agnostic context component contributes considerably more to rate reductions than the ITA layers that leverage inter-tile spatial dependencies. Still, we argue that the NTI-layers fulfill their purpose, to synergize with the partitioning strategy that maximizes processing throughput (Section 6.4).
6.3 Image Reconstruction Quality
We aim to demonstrate the feasibility of recovering presentable images from the compressed latent space of shallow features. We average results on DOTA-2, SpaceNet-3, and xView to reduce the bloat of reporting similar values summarize the results in Table V. For transparency, we exclusively select samples from the lower quartile across all measures to qualitatively showcase the reconstruction.
@@ 6.3 Image Reconstruction Quality
Model PSNR↑↑ MS-SSIM↑↑ LPIPS↓↓ BPP↓↓ Pred. Loss↓↓
FOOL 36.51 15.43 0.1480 0.1808 -
FOOL-FT 35.56 14.57 0.1700 0.1808 -
FP-HQ 43.22 25.07 0.0896 1.0470 1.500
@@
MSHP-HQ 43.90 25.20 0.0841 1.0370 1.711
MSHP-MQ 36.45 16.83 0.2361 0.2787 1.180
JAHP-HQ 43.95 25.17 0.0818 1.0297 1.504
@@
Figure 13 further visualizes the potential of reliably recovering fine-grained from the compressed latent space. Naturally, it should be possible to finetune the TinyLic-MQ to improve perceptual quality analogous to FOOl-FT. However, TinyLIC is still a significantly costlier model, with a worse rate and prediction performance. More pressingly, we stress that the reliability of a restoration model is bound by the available signals in the compressed latent space. Accordingly, we deliberately avoid generative models that prioritize realism over structural integrity. In particular, prioritizing realism over reliability defeats the primary purpose of image restoration, i.e., intervention by human experts in critical EO applications. A model outperforming experts does not imply that predictions may inexplicably be false. Specifically, where human cost is involved (e.g., disaster warning or relief [5]) it is paramount that experts can trust the codec to not include extrapolated elements to an image.
We argue that our results adequately underpin the statements in Section 3.3.1 and Section 3.3.2. In summary, if the salient regions align, compressing for model prediction requires more information than for human observation. Task specificity determines rate savings and not an entity’s input interface. Targeting shallow features is minimally task-specific by relaxing the objective for lossless prediction on all possible tasks to those valuable for clients.
@@ 6.4.1 Processing Throughput and Transfer Cost Reduction
Device Model Spatial Dim. Batch Size TCR/s ↑↑
FOOL-L 600x600 4 $7.4794\cdot 10^{8}$
@@
Orin Nano FOOL-S 600x600 8 $\boldsymbol{1.3386\cdot 10^{9}}$
FOOL-L 600x600 8 $7.5456\cdot 10^{8}$
@@
NX FOOL-S 600x600 8 $\boldsymbol{1.3422\cdot 10^{9}}$
@@
Orin Nano BSVBI-S 550x550 7 $6.9919\cdot 10^{9}$
@@ 6.4.2 Model Inference with Concurrent Task Execution
Model TCR/s [[[[conc]]]] TCR/s dec. File Size (MB) File/s rANS conc. (MB/s)
FOOL-L $7.26\cdot 10^{8}$ 2.94% 0.616 29 37.2
FOOL-M $7.57\cdot 10^{8}$ 1.28% 0.462 31 38.3
FOOL-S $1.32\cdot 10^{9}$ 1.06% 0.383 52 38.8
BSVBI-L $5.96\cdot 10^{8}$ 2.88% 0.822 37 37.6
BSVBI-M $7.15\cdot 10^{8}$ 1.37% 0.617 42 39.6
BSVBI-S $6.91\cdot 10^{8}$ 1.28% 0.437 58 40.1
@@
Figure 15: CPU (red) and GPU (blue) Usage of Encoder Network
@@
Figure 16: CPU (red) and GPU (blue) Usage of Concurrent Pipeline
@@ 6.4.3 Downlinkable Data Volume
Figure 17: Downlinkable Data Volumes by Link
Notice the logarithmic scale, i.e., FOOL improves downlinking using bent pipes by over two orders of magnitude without relying on prior information on the downstream tasks or crude filtering methods. For example, given Maxar’s WorldView-3 conditions [29], it would be possible to downlink roughly 9TB of sensor data per pass, before reaching downlink saturation. As a comparison, the state-of-the-art filtering method in [27] reports a 3$\times$ improvement, based on a definition of value. Note that to provide a realistic presentation of the opportunities SVBI provides, we assume that a nanosatellite processes tiles until reaching a downlink segment. Moreover, we disregard the “computational deadline”, i.e., it can process all the data before reaching a ground segment. This is reasonable since there should always be enough data to process. If not produced by a single onboard sensor, constellations may designate certain satellites as compression nodes using reliable, high-capacity local communication channels [78]. Further, from the throughput evaluation and the orbital period, it is inferrable that even the low-end current-generation Orin Nano would barely miss the computational deadline. We remind the reader that the presented throughput measures represent a worst-case lower bound, as no vendor-specific software or hardware optimization was applied.
@@ 6.4.4 Energy Consumption and Savings
Figure 18: Energy Cost of Compression Pipelines
As expected, processing on TX2 requires more energy than on NX and Orin Nano as it is slower. Somewhat interesting is that the NX consumes more energy than the Orin Nano. As they execute the same models with comparable processing throughput, the results suggest that the newer Jetson lineup is more energy-efficient. Lastly, we measure savings from reduced transmission time, arguably an often undervalued advantage of compression. Admittedly, satellites will downlink as bandwidth permits, i.e., the transmission energy cost does not depend on the codec performance when there is saturation. Nonetheless, to intuitively show the amount of energy large volumes might require, we contrast with bent pipes in Figure 19.
@@ 7 Discussions and Limitations
The following elaborates on limitations and highlights some research directions we find particularly promising to increase mission value. A general recommendation is to leverage recent advancements in image restoration [79] by utilizing generative models as robust priors and ensuring enough signals are available in the downlinked data to preserve structures at salient regions.